OpenStudy (anonymous):

The Cost of Cynthia's dinner is $15.20. She leaves a tip that is 15% of the cost of the dinner.What is the best estimate for the amount of the tip ?

OpenStudy (anonymous):

\(15+15\div 2=15+7.5=22.5\) estimate at \(\$2.25\)

OpenStudy (anonymous):

cheapskate

OpenStudy (anonymous):

2.28

OpenStudy (anonymous):

Thanks guys So It Is 2.00 Right ?

OpenStudy (anonymous):

i would say as an estimate \(\$2.25\) but @susan3 has it exactly

OpenStudy (anonymous):

Is \(15\%\) really cheap?

OpenStudy (anonymous):

Round it down to \($2\), lol